# JuStRank: Benchmarking LLM Judges for System Ranking

### v1

@@ 1 Introduction
Since relying on an inaccurate judge will likely result in sub-optimal decisions, this trend lends an urgency to evaluating the performance of the LLM judges themselves. Indeed, recent works attempt to benchmark judging capabilities, compiling leaderboards of judge performance Lambert et al. (2024); Tan et al. (2024) as well as analyzing their sensitivities and biases Wang et al. (2023); Thakur et al. (2024); Wei et al. (2024); Bavaresco et al. (2024); Feuer et al. (2024); Liu et al. (2024b); Lee et al. (2024a); Xu et al. (2024); Ye et al. (2024).
@@
In this work we aim to address this gap, and characterize the system-level evaluation capabilities and behaviors of LLM-based judges. To this end, we introduce a novel judge benchmark – JuStRank (Judges for System Ranking). JuStRank compares judges by their ability to correctly rank models, based on agreement with a ground-truth model ranking. JuStRank encompasses a collection of $48$ state-of-the-art judges, including both general-purpose LLMs and reward models. Our large-scale benchmark and analysis allow us to investigate the performance and behavior of judges when ranking systems.
@@ 2 The Gap in Judge Benchmarking
In this section, we outline why existing estimations of judge performance are insufficient to decide which judge is best at choosing between target systems. (Figure 1, bottom).
@@
For system-level judge evaluation, however, the error distribution plays a key role, as judge errors may distribute unevenly across systems, impacting their induced ranking. For example, a judge may exhibit an unjustifiable preference (positive bias) towards responses from a particular system $A$. Thus, this judge will tend to give this system the wrong ranking, even if it makes very few mistakes on responses from other systems (i.e., has an overall high instance-level accuracy). Hence, a more uniform distribution of errors – reflecting less biased judgment – is a desirable quality for system-level judges, and one that may lead to a more accurate ranking.
@@ 3 Task Formulation
Judges $\mathbf{J}=\{j_{p}\}_{p=1}^{P}$ map a pair of instruction $i_{k}$, and system response $r_{k}^{l}$ to a scalar score that estimates the quality of the response. Each judge has a specific realization for performing this score mapping [^2] We note that some realizations, such as the comparative realization in §4.2.2, may incorporate a separate set of responses to perform the judgment., of the form: $j_{p}(i_{k},r_{k}^{l})=Score_{k,l}^{p}$. Once a judge $j_{p}$ scores all $K\times L$ responses, we can define a scores matrix $j_{p}(R)\in\mathbb{R}^{K\times L}$ where $j_{p}(R)_{k,l}=Score_{k,l}^{p}$.
In order to quantify system-level quality, we must apply an aggregation method, $a\in A=\{a\colon\mathbb{R}^{K\times L}\longrightarrow\mathbb{R}^{L}\}$. The aggregation method $a$ maps a scores matrix $j_{p}(R)$ to a system-level vector $V^{p,a}\in\mathbb{R}^{L}$ where each entry, $V^{p,a}_{l}$, is a single overall quality score for system $s_{l}$ by judge $j_{p}$. In turn, ordering the systems scores in $V^{p,a}$ induces a ranking over the systems set $\mathbf{S}$.
@@ 4 Experimental setup
To explore judge performance and behavior, we utilize responses from multiple systems (§4.1) and run reward model judges (§4.2.1) and LLM judges (§4.2.2) over these responses. To obtain system rankings, we experiment with different aggregation methods (§4.3) over the judge scores. Finally, the resulting rankings are compared against a gold system ranking, obtained from a separate dataset (§4.4).
@@ 4.2.2 LLM Judge Realizations
The judge is asked to provide a quality score of the response on a Likert Likert (1932) scale with $5$ labels: [Very Bad, Bad, Mediocre, Good, Very Good]. We then convert the textual judgments to scores in $[1-5]$.
@@
In total, we collect judgments from $10$ LLMs and $4$ realizations [^3] Prompts for all realizations are provided in Appendix G., yielding $40$ LLM judges. We use the following generative LLM judges: Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Mixtral-8x22B-Instruct-v0.1, Mistral-Large-Instruct-2407, Llama-3.1-405B-Instruct Dubey et al. (2024), Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct, Qwen2.5-72B-Instruct, GPT-4o and GPT-4o-mini.
@@ 6.1 Some Judges Are Particularly Decisive
Figure 5 depicts the relationship between predicted win-rates and gold win-rates for several judges. The quadrants in the figure indicate whether the judge’s pairwise preference decision is aligned with the gold preference. As can be expected, the judge predictions in Figure 5 are often centered around the ground-truth win-rates determined by humans. But strikingly, some judges exhibit unique prediction patterns, yielding win-rates that are consistently closer to the extremes ($0.0$ / $1.0$) compared to the human data. For instance, for pairs with a ground-truth win-rate of $\sim$0.80.80.80.8, we can see that the predicted win-rate in the judgments of Llama-405B (Fig. 5, right) tends to exceed $0.9$. Put simply, when faced with a response from a strong system, the judge is very likely to prefer it over the response of a less capable system, even where human judges are less decisive.
This sigmoidal win-rate prediction pattern resembles behaviors previously described for classifier calibration Silva Filho et al. (2023), where classifiers may exhibit “overconfidence” in their predicted probabilities. [^5] Note, however, that the behavior in our case does not reflect judge probability scores, but rather the empirical ratio of instances where the responses $\{r_{k}^{l}\}_{l=1}^{l=L}$ of a system $k$ are preferred over those of another system. Thus, following Kull et al. (2017), we quantify judges’ decisive (overconfident) behavior by fitting the cumulative beta distribution function to the win-rate prediction plots. This enables describing judge prediction behavior in terms of a single fit value $\alpha=\beta$, where $\alpha\in[0,\infty]$, $\alpha=1$ represents no over- or under-decisiveness, and larger values represent more decisive behavior (refer to Appendix F for details). Figure 6a and App. Fig. 11 depict the beta curve fit for win-rates of various judges.
@@ 6.2 Bias Towards Specific Systems
A major concern when using judges for system preference is judge bias – a judge may treat a specific system “unfairly”, by consistently judging its responses too favorably or too harshly.
We define the bias $B_{s_{a}}^{p}$ of judge $j_{p}$ towards system $s_{a}$ by the expectation over the differences between the predicted win-rate and the gold win-rate, over all systems that $s_{a}$ interacts with. Formally, $B_{s_{a}}^{p}=\mathbb{E}_{s_{b}\in S}(WR^{p}(s_{a},s_{b})-WR^{g}(s_{a},s_{b}))$. In other words, if according to $j_{p}$ the win-rates of system $s_{a}$ are (on average) higher than those in the human data, we will say that $j_{p}$ exhibits positive bias towards it; and if they are lower than the ground-truth, $j_{p}$ would be said to exhibit negative bias towards it.
@@
We also ask whether LLM judges exhibit self-bias Xu et al. (2024), i.e., bias towards the system that uses the same underlying LLM. While we find some instances of self-bias, this is not a consistent effect across judge realizations (App. Table 3).
@@ 7 Related Work
For LLM-based judges and reward models, prior works have opted for an instance-level evaluation approach, curating benchmarks of task outputs with ground-truth quality annotations in order to evaluate judge performance. Most prominently, RewardBench Lambert et al. (2024) compares dozens of judges (including reward models, generative LLMs, and classifiers) on the task of correctly deciding between pairs of outputs, labeled as "preferred" or "rejected" by human annotators. RewardBench aims to identify the most suitable judges for model alignment, e.g., for use in RLHF; in contrast, the present work measures judges in terms of their ability to compare the performance of candidate systems. Another recent instance-level benchmark is JudgeBench Tan et al. (2024), which focuses on curating challenging response pairs where the judge must discern subtle errors.
@@
Several popular benchmarks rely on LLM judges to produce leaderboards of state-of-the-art systems. Such benchmarks – e.g., Arena Hard Li et al. (2024) and AlpacaEval Dubois et al. (2024) – do perform a system-level validation of their resulting leaderboards against other benchmark rankings (see Perlitz et al., 2024). However, such efforts are limited to validating the particular dataset and judge setup chosen for the benchmark (usually incorporating GPT-4 as the judge), rather than comparing and analyzing the performance of different judge models and implementations. Thakur et al., 2024 conduct a task-specific system-level evaluation of judges, over the TriviaQA Joshi et al. (2017) dataset. Compared to their work, the present study is on a larger scale and offers novel metrics and analyses on system-level judge behaviors.
@@ 8 Discussion
Our system-level approach has multiple additional benefits. First, it forces the evaluation of judges to be representative with respect to the distribution of systems that generate the responses. In existing instance-level benchmarks this factor is not taken into account, and likely results in less accurate judge evaluations.
Second, it affords a new perspective on what it means for a judge to be biased; on the one hand, we discover some decisiveness trends (§6.1) that may actually be useful for making correct preference decisions, and increasing the separability between systems; and on the other, we report some problematic biases that directly distort the judgment of particular systems (§6.2). An important avenue for future work is to connect our findings here to the existing literature on judge biases Ye et al. (2024), and understand to what extent both of these behaviors stem from particular LLM style attributes Feuer et al. (2024).
Given this vast and complex space, our work is admittedly only a first step in understanding the behavior of judges for ranking and selecting LLMs. We encourage the community to explore these issues further: for instance, by training dedicated system-level judges, exploring judge ensembles, or studying other aggregation approaches. We believe that JuStRank can facilitate such research directions, as it can be easily extended to new judges without requiring additional human annotations.
Our hope is that both practitioners and researchers can benefit from JuStRank, by making more informed choices of judges to suit their needs.
@@ 9 Conclusion
In this work we conducted the first comprehensive evaluation of system ranking by LLM judges. We tested a wide array of judges, including reward models, as well as different realizations of generative LLMs, over a large collection of systems. We collected system responses over a diverse set of instructions. The judges scored each response, and we compiled a ranking by aggregating the judgments over all the responses. Then, the quality of the judge’s system ranking was compared to a human-based ranking, producing the JuStRank leaderboard.
@@
Our analyses of LLM judge realizations are, by necessity, limited to the specific realization prompts that we used. Several studies show that LLMs Mizrahi et al. (2024) as well as LLM judges Wei et al. (2024) are brittle with respect to prompt phrasing, and hence this may have had an impact on the results.
@@ Appendix C Chatbot Arena Data
We use their public data file from August 2024777Chatbot Arena data, and follow the official notebook00footnotemark: 0 to extract the raw data, deduplicate it, and calculate the overall system rankings. This dataset includes the human preference judgments and names of the participating systems, but not the instructions or system responses for the battles.

### v2

@@ 1 Introduction
Since relying on an inaccurate judge will likely result in sub-optimal decisions, this trend lends an urgency to evaluating the performance of the LLM judges themselves. Indeed, recent works attempt to benchmark judging capabilities, compiling leaderboards of judge performance Lambert et al. (2024); Tan et al. (2024) as well as analyzing their sensitivities and biases Wang et al. (2023); Wei et al. (2024); Bavaresco et al. (2024); Feuer et al. (2024); Liu et al. (2024b); Xu et al. (2024); Ye et al. (2024).
@@
In this work we aim to address this gap, and characterize the system-level evaluation capabilities and behaviors of LLM-based judges. To this end, we introduce a novel judge benchmark – JuStRank (Judges for System Ranking). JuStRank compares judges by their ability to correctly rank models, based on agreement with a ground-truth model ranking. JuStRank encompasses a collection of $48$ state-of-the-art judges, including both general-purpose LLMs and reward models. Our large-scale benchmark and analysis allow us to explore the performance and behavior of judges as system rankers.
@@
4. To facilitate further research into judge behavior, we release our data [^1] JuStRank Judge Scores data, comprising $1.5$M judgment scores given by LLMs and reward models.
@@ 2 The Gap in Judge Benchmarking
In this section, we outline why existing estimations of judge performance are insufficient to decide which judge is best at choosing between target systems (Figure 1, bottom).
@@
For system-level judge evaluation, however, the error distribution plays a key role, as judge errors may distribute unevenly across systems, impacting their induced ranking Dorner et al. (2024); von Däniken et al. (2024). For example, a judge may exhibit an unjustifiable preference (positive bias) for responses from a particular system $A$. Thus, this judge will tend to give $A$ an incorrect ranking, even if it makes very few mistakes on responses from other systems (i.e., has an overall high instance-level accuracy). Hence, a more uniform distribution of errors – reflecting less biased judgment – is a desirable quality for system-level judges, and one that may lead to a more accurate ranking.
@@ 3 Task Formulation
Judges $\mathbf{J}=\{j_{p}\}_{p=1}^{P}$ map a pair of instruction $i_{k}$, and system response $r_{k}^{l}$ to a scalar score that estimates the quality of the response. Each judge has a specific realization for performing this score mapping333 We note that some realizations, such as the comparative realization in §4.2.2, may incorporate a separate set of responses to perform the judgment., of the form: $j_{p}(i_{k},r_{k}^{l})=Score_{k,l}^{p}$. Once a judge $j_{p}$ scores all $K\times L$ responses, we can define a scores matrix $j_{p}(R)\in\mathbb{R}^{K\times L}$ where $j_{p}(R)_{k,l}=Score_{k,l}^{p}$.
In order to quantify system-level quality, we must apply an aggregation method, $a\in A=\{a\colon\mathbb{R}^{K\times L}\longrightarrow\mathbb{R}^{L}\}$. The aggregation method $a$ maps a scores matrix $j_{p}(R)$ to a system-level vector $V^{p,a}\in\mathbb{R}^{L}$ where each entry, $V^{p,a}_{l}$, is a single overall quality score for system $s_{l}$ by judge $j_{p}$. In turn, ordering the system scores in $V^{p,a}$ induces a ranking over the systems set $\mathbf{S}$.
@@ 4 Experimental setup
To explore judge performance and behavior, we utilize responses from multiple systems (§4.1) and run reward model judges (§4.2.1) and LLM judges (§4.2.2) over these responses. To obtain system rankings, we experiment with different aggregation methods (§4.3) over the judge scores. Finally, the resulting rankings are compared against a gold system ranking, taken from a separate dataset (§4.4).
@@ 4.2.2 LLM Judge Realizations
The judge provides a quality score of the response on a Likert Likert (1932) scale with $5$ labels: [Very Bad, Bad, Mediocre, Good, Very Good]. We then convert the textual judgments to scores in $[1-5]$.
@@
In total, we collect judgments from $10$ LLMs and $4$ realizations, yielding $40$ LLM judges. Prompts for all realizations are provided in Appendix G.
We use the following generative LLM judges: Llama-3.1-405B-Instruct Dubey et al. (2024), Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3-70B-Instruct, Mixtral-8x22B-Instruct-v0.1, Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Mistral-Large-Instruct-2407, Qwen2.5-72B-Instruct, GPT-4o and GPT-4o-mini.
@@ 6.1 Some Judges are Particularly Decisive
[…]
This sigmoidal win-rate prediction pattern resembles behaviors previously described for classifier calibration Silva Filho et al. (2023), where classifiers may exhibit “overconfidence” in their predicted probabilities. [^5] Note, however, that the behavior in our case does not reflect judge probability scores, but rather the empirical ratio of instances where the responses $\{r_{k}^{l}\}_{l=1}^{l=L}$ of a system $k$ are preferred over those of another system. Thus, following Kull et al. (2017), we quantify judges’ decisive (overconfident) behavior by fitting the cumulative beta distribution function to the win-rate prediction plots. This enables describing judge prediction behavior in terms of a single fit value $\alpha=\beta$, where $\alpha\in[0,\infty]$, a value of $\alpha=1$ represents no over- or under-decisiveness, and larger values represent more decisive behavior (refer to Appendix F for details). Figure 6a and App. Fig. 11 depict the beta curve fit for win-rates of various judges.
@@ 6.2 Bias Towards Specific Systems
A major concern when using judges for system preference is judge bias – a judge may treat a specific system “unfairly”, by consistently judging its responses too favorably or too harshly (see Von Däniken et al., 2024).
We define the bias $B_{s_{a}}^{p}$ of judge $j_{p}$ towards system $s_{a}$ by the expectation over the differences between the predicted and gold win-rates, over all systems that $s_{a}$ interacts with. Formally, $B_{s_{a}}^{p}=\mathbb{E}_{s_{b}\in S}(WR^{p}(s_{a},s_{b})-WR^{g}(s_{a},s_{b}))$. [^6] Our formulation of bias aims to reflects the practical impact of the judge bias on system preference. This is in contrast to the Favi-Score metric proposed by Von Däniken et al. (2024), which is decoupled from the overall accuracy of preference decisions. In other words, if according to $j_{p}$ the win-rates of system $s_{a}$ are (on average) higher than those in the human data, we will say that $j_{p}$ exhibits positive bias towards it; and if they are lower than the ground-truth, $j_{p}$ would be said to exhibit negative bias.
@@
We also ask whether LLM judges exhibit self-bias Xu et al. (2024); Panickssery et al. (2024), i.e., bias towards the system that uses the same underlying LLM. While we find some instances of self-bias, this is not a consistent effect across judge realizations (App. Table 3).
@@ 7 Related Work
Prior works on LLM-based judges have opted for an instance-level evaluation approach, curating benchmarks of responses with ground-truth quality annotations in order to evaluate judge performance. Most prominently, RewardBench Lambert et al. (2024) compares dozens of judges (including reward models, generative LLMs, and classifiers) on the task of deciding between pairs of outputs. RewardBench aims to identify the most suitable judges for model alignment, e.g., for use in RLHF; in contrast, our work measures judges in terms of their ability to compare the performance of candidate systems. Another recent instance-level benchmark, JudgeBench Tan et al. (2024), focuses on curating challenging response pairs where the judge must discern subtle errors.
@@
Several popular benchmarks rely on LLM judges to produce leaderboards of state-of-the-art systems. Such benchmarks – e.g., Arena Hard Li et al. (2024) and AlpacaEval Dubois et al. (2024) – do perform a system-level validation of their resulting leaderboards against other benchmark rankings (see Perlitz et al., 2024). However, such efforts are limited to validating the particular dataset and judge setup chosen for the benchmark (usually with GPT-4 as the judge), rather than comparing and analyzing the performance of different judge models and implementations. Thakur et al., 2024 conduct a task-specific system-level evaluation of judges, over the TriviaQA Joshi et al. (2017) dataset. Compared to their work, the present study is on a larger scale and offers novel metrics and analyses on system-level judge behaviors.
@@ 8 Discussion
Our system-level approach has multiple additional benefits. First, it forces the evaluation of judges to be representative with respect to the distribution of systems that generate the responses. In existing instance-level benchmarks this factor is not taken into account, and likely results in less accurate judge evaluations. Second, it affords a new perspective on what it means for a judge to be biased; on the one hand, we discover some decisiveness trends (§6.1) that may actually be useful for making correct preference decisions, and increasing the separability between systems; and on the other, we report some problematic biases that directly distort the judgment of particular systems (§6.2). An important avenue for future work is to connect our findings here to the existing literature on judge biases Ye et al. (2024), and understand to what extent both of these behaviors stem from particular LLM style attributes Feuer et al. (2024).
Given this vast and complex space, our work is admittedly only a first step in understanding the behavior of judges for ranking and selecting LLMs. We release our raw judgment scores data, and encourage the community to explore these issues further: for instance, by training dedicated system-level judges, exploring judge ensembles, or studying other aggregation approaches. We believe that JuStRank can facilitate such research directions, as it can be easily extended to new judges without requiring additional human annotations.
Our hope is that both practitioners and researchers can benefit from JuStRank, by making more informed choices of judges for their needs.
@@ 9 Conclusion
In this work we conducted the first comprehensive evaluation of system ranking by LLM judges. We tested a wide array of judges, including reward models and different realizations of generative LLMs, over a large collection of systems.
We collected system responses over a diverse set of instructions. The judges scored each response, and we compiled a ranking by aggregating the judgments over all responses. Then, the quality of the judge’s system ranking was compared to a human ranking, producing the JuStRank leaderboard.
@@
Our analyses of LLM judge realizations are, by necessity, limited to the specific realization prompts that we used. Several studies show that LLMs Mizrahi et al. (2024) as well as LLM judges Wei et al. (2024) are brittle with respect to prompt phrasing, and hence this may have had an impact on the results. In addition, there can be some variations in judge responses depending on the exact API and inference implementation used.
@@ Appendix C Chatbot Arena Data
We use their public data file from August 2024888Chatbot Arena data, and follow the official notebook00footnotemark: 0 to extract the raw data, deduplicate it, and calculate the overall system rankings. This dataset includes the human preference judgments and names of the participating systems, but not the instructions or system responses for the battles.